# AKRMap: Adaptive Kernel Regression for Trustworthy Visualization of Cross-Modal Embeddings

### v1

@@ 1 Introduction
Cross-modal embeddings play a fundamental role for multi-modal models, functioning as cross-modal encoders (Rombach et al., 2022), objective functions (Ramesh et al., 2022), or evaluation metrics (Hessel et al., 2021; Wu et al., 2023a) for tasks like text-to-image (T2I) generation. To assess the alignment of multi-modal models, various evaluation metrics based on these embeddings have been introduced, such as CLIPScore (Hessel et al., 2021), and Human Preference Score (HPS) (Wu et al., 2023a, b; Zhang et al., 2024). Despite their utility, embedding-based evaluations are often difficult to interpret, as the metrics are typically reported as aggregated values, without providing insights into instance-level performance. This deficiency undermines the trustworthiness and transparency of the evaluation.
To address this limitation, dimensionality reduction (DR) visualization serves as a crucial tool, offering a way to reveal the landscape of cross-modal metrics and enabling a more comprehensive understanding of model performance. Recent studies (Liang et al., 2022; Wang et al., 2023b, c) have explored the use of DR methods for cross-modal embeddings. These efforts often leverage established techniques such as PCA (Li et al., 2018), UMAP (McInnes et al., 2018) and t-SNE (Van der Maaten & Hinton, 2008), as well as autoencoder-based approaches (Le et al., 2018; Elhamod & Karpatne, 2024). However, existing DR methods are primarily designed to depict feature distributions within a single modality. When applied to cross-modal metrics, these methods often produce dense neighborhoods where points with significantly different metric values are positioned close, leading to overlaps and local occlusion (Figure 1(b)). Moreover, multi-modal models are typically evaluated on large-scale datasets containing millions of data points. This creates a need for rendering contour maps that can reveal continuous metric distributions. However, local neighboring points with a mix of high and low metric values may be misrepresented, as the contour map depicts only a single aggregated value, causing inaccurate contour mapping as illustrated in Figure 1(c).
To construct trustworthy visualizations for cross-modal embeddings, a key consideration is to enhance the accuracy of metric contour mapping in the projected 2D space. As contour estimation typically relies on radius basis function (RBF) kernel (e.g., Gaussian kernel), we seek to enhance the kernel-based mapping by coordinating it with the DR process. Drawing inspiration from supervised t-SNE (Gisbrecht et al., 2015) and UMAP (Sainburg et al., 2021), we introduce Adaptive Kernel Regression Map (AKRMap), a supervised DR method that leverages adaptive kernel regression to effectively visualize the distribution of cross-modal embedding metrics. Specifically, AKRMap first constructs a supervised DR network explicitly guided by post-projection kernel regression loss with neighborhood constraint (Sect. 4.1.1). Next, to account for the gap between high-dimensional and low dimensional kernels, we improve the flexibility of the post-projection kernel through adaptive generalized kernel jointly learned with the projection (Sect. 4.1.2). This approach enables the generation of scatterplots for discrete data points and contour maps for continuous metric landscapes, while also supporting advanced features such as overlay views and zooming for interactive interpretation (Sect. 4.2). Both the DR method and the visualization tool have been implemented as a Python package and will be made publicly available.
We conduct quantitative experiments to evaluate AKRMap against traditional and autoencoder-based DR visualizations (Sect. 5.1). The results highlight the superior performance of AKRMap in accurately mapping cross-modal metrics for both in-sample and out-of-sample data points. We further demonstrate the practical applications of AKRMap across three distinct scenarios (Sect. 5.2): 1) visual exploration of human preference dataset (HPD) (Wu et al., 2023a), 2) visual comparison of diffusion-based and auto-regressive T2I models, and 3) visual examination of the global impact for fine-tuning. These applications illustrate that AKRMap effectively enables human-in-the-loop interpretation of model performance on large-scale datasets.
@@
  • We develop a tool for trustworthy visualization of cross-modal metrics, incorporating visualization features such as a scatterplot view and contour maps, along with interactive features like zooming and overlaying.
  • We conduct quantitative experiments to demonstrate the superior performance of AKRMap in generating more accurate visualizations of cross-modal metric, and highlight its applications across three scenarios to enhance the trustworthiness of T2I model evaluations.
@@ 2.1 Cross-Modal Embedding-based Evaluation
Cross-modal embeddings like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) form the foundation of multi-modal learning. Specifically, many evaluation methods for multi-modal models, such as T2I models, rely on these embeddings. Conventional metrics like Inception Score (Salimans et al., 2016) and Fréchet inception distance (Heusel et al., 2017) compute the average distance or distributional difference between generated and reference images in the embedding space, yet they are insufficient to measure cross-modal alignment and instance-level performance. With the advancement of multi-modal AI, cross-modal embedding metrics like CLIPScore (Hessel et al., 2021) have emerged to measure the alignment between prompts and generated images. While being effective in capturing the alignment, CLIPScore fails to model human preferences. To address this limitation, recent studies have proposed training specialized human preference models such as HPS (Wu et al., 2023a, b; Zhang et al., 2024), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2024), to better align with human judgments. For example, ImageReward (Xu et al., 2024) introduces a systematic pipeline for expert annotations to train preference models, while PickScore (Kirstain et al., 2023) leverages large-scale crowd-sourced human comparisons.
@@
Figure 2: AKRMap is a neural network based DR method designed to learn adaptive kernel regression for visualizing cross-modal embeddings. The network integrates two key components to jointly learn data point projections and cross-modal metric estimation: 1) Kernel regression supervision, and 2) Adaptive generalized kernel. The resulting visualizations, including scatterplots and contour maps, provide a clearer and more accurate representation of the cross-modal metric distribution.
@@ 2.2 Visualization for Neural Embeddings
Visualization is an important tool to support data analysis in conjunction with AI (Wang et al., 2023a; Ye et al., 2024). Particularly, it has proven to be effective for enhancing the understanding of various types of neural embeddings, including word embeddings (Mikolov et al., 2013; Chiang et al., 2020), vision-language embeddings (Liang et al., 2022; Wang et al., 2023b; Ye et al., 2025), and parameter spaces within loss landscape (Li et al., 2018; Elhamod & Karpatne, 2024). Commonly, dimensionality reduction (DR) techniques are employed for visualization. These include linear DR methods such as PCA (Abdi & Williams, 2010), as well as nonlinear DR methods like t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018). Parametric versions of traditional DR methods, which provide explicit mapping for projections, can be achieved by training neural networks, such as parametric t-SNE (Gisbrecht et al., 2015; Damrich et al., 2023), parametric UMAP (Sainburg et al., 2021) and other parametric methods for interactive clustering (Xia et al., 2023) or streaming data (Xia et al., 2024). Similarly, auto-encoder based DR methods (Le et al., 2018; Elhamod & Karpatne, 2024) have also been developed, incorporating an additional objective of reconstructing high-dimensional embeddings from their projections. Recently, there has been a growing interest in utilizing DR techniques to visualize cross-modal embeddings. For instance, UMAP has been employed to explore and visualize the modality gap between text and image embeddings (Liang et al., 2022; Wang et al., 2023b). To address this modality gap and enhance the cross-modal alignment in visualization space, some studies have introduced fusion-based DR methods (Ye et al., 2025), enabling the visualization of image embeddings in relation to text embedding anchor points.
@@
However, previous efforts treat multi-modal embeddings separately, failing to effectively depict the landscape of cross-modal metrics (e.g., CLIPScore and HPS). In addition, existing metric-aware DR techniques focus on a limited range of particular metrics like distance (Sainburg et al., 2021) and density (Narayan et al., 2021), without a generalizable mechanism to preserve other different metrics. To address this limitation, we propose a kernel regression-based supervised projection method combined with an adaptive generalized kernel. Different from traditional kernel DR methods like kernel PCA (Schölkopf et al., 1997) that mainly use kernel to transform the high-dimensional features, we develop an adaptive kernel in the projection space to guide the DR. Notably, to construct the kernel guidance, we propose a novel cross-validation supervision trick that bridges between traditionally nonparametric kernel regression and parametric DR. This design allows the propagation of contour estimation errors back into the projection process. In this way, our method can dynamically adjust both the DR mapping and the kernel shape, enabling it to accurately model and fit the complex landscape of cross-modal metrics.
@@ 3 Problem Definition
Next, to visualize a continuous distribution from the discrete projected sample points in $P(e(D))$, a contour map in the 2D space needs to be estimated. Here, we adopt the approach used by contouring algorithms in various Python libraries, such as Plotly, by dividing the projected 2D space into a grid and calculating the metric distribution values at each grid point. Specifically, suppose the projected coordinates lie within a normalized 2D space of $[0,1]\times[0,1]$. For each grid $\mathbb{x}_{g}=(\frac{i}{N_{w}},\frac{j}{N_{h}})$, we compute a value $\hat{s}(\mathbb{x}_{g})$ based on the local metric distribution, which is then colored using a continuous colormap. $N_{w}$, $N_{h}$ denotes the number of grids along the x-axis and y-axis, respectively. In this manner, we can generate a contour map depicting continuous landscape of the metric distribution from the discrete projected sample points.
@@ 4.1.1 Kernel Regression Supervised Projection
where the overall regression loss $MSE_{r}$ is a weighted sum of the loss on validation set ($MSE_{vl}$) and that on training set ($MSE_{tr}$). We seek to balance these loss terms to ensure mapping accuracy at both in-sample and out-of-sample positions on the contour, and the weights $w_{1}=1$ and $w_{2}=0.3$ are set empirically.
@@ 4.2 Cross-modal Metric Visualization
  • Overlay. The scatterplot view and contour map allow data points to be overlaid onto the contour map, enabling users to simultaneously observe the overall distribution while interacting with individual data points, as shown in Figure 4. To mitigate occlusion, we have implemented sampling techniques, including random sampling and Poisson disk sampling.
@@ 5 Experiment
Table 1: Quantitative comparison of cross-modal embedding metric visualizations on the HPD dataset. Our AKRMap outperforms baseline methods and ablations, achieving the best performance across all four cross-modal embedding metrics.
@@ 5.1 Quantitative Experiments
Baseline methods. We compare AKRMap against three commonly used DR methods: PCA, t-SNE and UMAP, as well as two encoder-based approaches: SAE (Le et al., 2018) and Neuro-Visualizer (Elhamod & Karpatne, 2024). For the encoder-based methods, the decoder is used to estimate grid colors, while the other baseline methods rely on the common Gaussian RBF kernel to estimate the 2D distribution map. We use distance threshold to cut off empty areas for traditional DR techniques but keep the full landscape of autoencoder methods because far away positions in traditional DR results are ambiguous and cannot hardly be mapped back to original high-dimensional space. Additional details are provided in Appendix C.
@@
Cross-modal Embedding Metrics. We compare AKRMap against baseline methods for visualizing several commonly used cross-modal embedding metrics in T2I generation, including CLIPScore (Hessel et al., 2021), HPSv2 (Wu et al., 2023a), PickScore (Kirstain et al., 2023), and Aesthetic Score (Schuhmann et al., 2022).
@@ 5.2.1 Visualizing Large T2I Dataset
In alignment with the results from the quantitative experiments, our method demonstrates superior performance by accurately depicting score distributions with rich local details. In contrast, the baselines suffer from various limitations, such as over-smoothing due to averaging effects, which obscure local structures. For example, PCA introduces pronounced block effects with sharp edges at region boundaries, significantly disrupting the overall smoothness of the visualizations. Similarly, UMAP and t-SNE often struggle with low-value regions being either overshadowed by high values or averaged out, making it challenging to identify clusters with suboptimal performance. Surprisingly, Neuro-Visualizer performs the worst among the baselines for visualizing cross-modal embedding metrics. Except the Aesthetic Score, its results are highly unstable, with excessively large estimations in out-of-sample areas, and the contour maps are riddled with jagged terrains, exhibiting poor smoothness. This further underscores the strength of our AKRMap in producing more accurate and visually coherent representations. Overall, our AKRMap achieves a superior balance between local accuracy and smoothness. Furthermore, the mapping performance of traditional methods can deteriorate significantly if a smaller bandwidth is manually set, as demonstrated by additional results in Appendix C.

### v2

@@ 1 Introduction
Cross-modal embeddings play a fundamental role for multi-modal models, functioning as cross-modal encoders (Rombach et al., 2022), objective functions (Ramesh et al., 2022), or evaluation metrics (Hessel et al., 2021; Wu et al., 2023a) for tasks like text-to-image (T2I) generation. To assess the alignment of multi-modal models, various evaluation metrics based on these embeddings have been introduced, such as CLIPScore (Hessel et al., 2021) and Human Preference Score (HPS) (Wu et al., 2023a, b; Zhang et al., 2024). Despite their utility, embedding-based evaluations are often difficult to interpret, as the metrics are typically reported as aggregated values, without providing insights into instance-level performance. This deficiency undermines the trustworthiness and transparency of the evaluation.
To address this limitation, dimensionality reduction (DR) visualization serves as a crucial tool, offering a way to reveal the landscape of cross-modal metrics and enabling a more comprehensive understanding of model performance. Recent studies (Liang et al., 2022; Wang et al., 2023b, c) have explored the use of DR methods for cross-modal embeddings. These efforts often leverage established techniques such as PCA (Li et al., 2018), UMAP (McInnes et al., 2018) and t-SNE (Van der Maaten & Hinton, 2008), as well as autoencoder-based approaches (Le et al., 2018; Elhamod & Karpatne, 2024). However, existing DR methods are primarily designed to depict feature distributions within a single modality. When applied to cross-modal metrics, these methods often produce dense neighborhoods where points with significantly different metric values are positioned close, leading to overlap and local occlusion (Figure 1(b)). Moreover, multi-modal models are typically evaluated on large-scale datasets containing millions of data points. This creates a need for rendering contour maps that can reveal continuous metric distributions. However, local neighboring points with a mix of high and low metric values may be misrepresented, as the contour map depicts only a single aggregated value, causing inaccurate contour mapping as illustrated in Figure 1(c).
To construct trustworthy visualizations for cross-modal embeddings, a key consideration is to enhance the accuracy of metric contour mapping in the projected 2D space. As contour estimation typically relies on radial basis function (RBF) kernel (e.g., Gaussian kernel), we seek to enhance the kernel-based mapping by coordinating it with the DR process. Drawing inspiration from supervised t-SNE (Hajderanj et al., 2019), we introduce Adaptive Kernel Regression Map (AKRMap), a supervised DR method that leverages adaptive kernel regression to effectively visualize the distribution of cross-modal embedding metrics. Specifically, AKRMap first constructs a supervised DR network explicitly guided by post-projection kernel regression loss with neighborhood constraint (Sect. 4.1.1). Next, to account for the gap between high-dimensional and low dimensional kernels, we improve the flexibility of the post-projection kernel through adaptive generalized kernel jointly learned with the projection (Sect. 4.1.2). This approach enables the generation of scatterplots for discrete data points and contour maps for continuous metric landscapes, while also supporting advanced features such as overlay views and zooming for interactive interpretation (Sect. 4.2). Both the DR method and the visualization tool have been implemented as a Python package and made publicly available.
@@
We conduct quantitative experiments to evaluate AKRMap against traditional and autoencoder-based DR visualizations (Sect. 5.1). The results highlight the superior performance of AKRMap in accurately mapping cross-modal metrics for both in-sample and out-of-sample data points. We further demonstrate the practical applications of AKRMap across three distinct scenarios (Sect. 5.2): 1) visual exploration of human preference dataset (HPD) (Wu et al., 2023a), 2) visual comparison of diffusion-based and auto-regressive T2I models, and 3) visual examination of the global impact for fine-tuning. These applications illustrate that AKRMap effectively enables human-in-the-loop interpretation of model performance on large-scale datasets. Our method can also be potentially generalized to other modalities such as text-to-video task, with additional quantitative experiments shown in Appendix F.
@@
  • We develop a tool for trustworthy visualization of cross-modal metrics, incorporating visualization features such as a scatterplot view and a contour map, along with interactive features like zooming and overlaying.
  • We conduct quantitative experiments to demonstrate the superior performance of AKRMap in generating more accurate visualizations of cross-modal metric, and highlight its applications across three scenarios to enhance the trustworthiness of T2I model evaluation.
@@ 2.1 Cross-Modal Embedding-based Evaluation
Cross-modal embeddings like CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) form the foundation of multi-modal learning. Specifically, many evaluation methods for multi-modal models, such as T2I models, rely on these embeddings. Conventional metrics like Inception Score (Salimans et al., 2016) and Fréchet Inception Distance (Heusel et al., 2017) compute the average distance or distributional difference between generated and reference images in the embedding space, yet they are insufficient to measure cross-modal alignment and instance-level performance. With the advancement of multi-modal AI, cross-modal embedding metrics like CLIPScore (Hessel et al., 2021) have emerged to measure the alignment between prompts and generated images. While being effective in capturing the alignment, CLIPScore fails to model human preferences. To address this limitation, recent studies have proposed training specialized human preference models such as HPS (Wu et al., 2023a, b; Zhang et al., 2024), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2024), to better align with human judgments. For example, ImageReward (Xu et al., 2024) introduces a systematic pipeline for expert annotations to train preference models, while PickScore (Kirstain et al., 2023) leverages crowd-sourced human comparisons.
@@
Figure 2: AKRMap is a neural network based DR method designed to learn adaptive kernel regression for visualizing cross-modal embeddings. The network integrates two key components to jointly learn data point projection and cross-modal metric estimation: 1) Kernel regression supervision, and 2) Adaptive generalized kernel. The resulting visualizations, including scatterplots and contour maps, provide a clearer and more accurate representation of the cross-modal metric distribution.
@@ 2.2 Visualization for Neural Embeddings
Visualization is an important tool to support data analysis in conjunction with AI (Wang et al., 2023a; Ye et al., 2024). Particularly, it has proven to be effective for enhancing the understanding of various types of neural embeddings, including word embeddings (Mikolov et al., 2013; Chiang et al., 2020), vision-language embeddings (Liang et al., 2022; Wang et al., 2023b; Ye et al., 2025), and parameter spaces within loss landscape (Li et al., 2018; Elhamod & Karpatne, 2024). Commonly, dimensionality reduction (DR) techniques are employed for visualization. These include linear DR methods such as PCA (Abdi & Williams, 2010), as well as nonlinear DR methods like t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018). Parametric versions of traditional DR methods, which provide explicit mapping for projections, can be achieved by training neural networks, such as parametric t-SNE (Gisbrecht et al., 2015; Damrich et al., 2023), parametric UMAP (Sainburg et al., 2021) and other parametric methods for tasks like interactive clustering (Xia et al., 2023) or streaming data (Xia et al., 2024). Similarly, autoencoder-based DR methods (Le et al., 2018; Elhamod & Karpatne, 2024) have also been developed, incorporating an additional objective of reconstructing high-dimensional embeddings from their projections. Recently, there has been a growing interest in utilizing DR techniques to visualize cross-modal embeddings. For instance, UMAP has been employed to explore and visualize the modality gap between text and image embeddings (Liang et al., 2022; Wang et al., 2023b). To address this modality gap and enhance the cross-modal alignment in visualization space, some studies have introduced fusion-based DR methods (Ye et al., 2025), enabling the visualization of image embeddings in relation to text embedding anchor points.
@@
However, previous efforts treat multi-modal embeddings separately, failing to effectively depict the landscape of cross-modal metrics (e.g., CLIPScore and HPS). In addition, existing metric-aware DR techniques focus on a limited range of particular metrics like distance (Sainburg et al., 2021) and density (Narayan et al., 2021), without a generalizable mechanism to preserve other different metrics. To address this limitation, we propose a kernel regression-based supervised projection method combined with an adaptive generalized kernel. Different from traditional kernel DR methods like kernel PCA (Schölkopf et al., 1997) that mainly use kernel to transform the high-dimensional features, we develop an adaptive kernel in the projection space to guide the DR. Notably, to construct the kernel guidance, we propose a novel cross-validation supervision technique that bridges between traditionally nonparametric kernel regression and parametric DR. This design allows the propagation of contour estimation errors back into the projection process. In this way, our method can dynamically adjust both the DR mapping and the kernel shape, enabling it to accurately fit the complex landscape of cross-modal metrics.
@@ 3 Problem Definition
Next, to visualize a continuous distribution from the discrete projected sample points in $P(e(D))$, a contour map in the 2D space needs to be estimated. Here, we adopt the approach used by contouring algorithms in various Python libraries, such as Plotly, by dividing the projected 2D space into a grid and calculating the metric distribution values at each grid point. Specifically, suppose the projected coordinates lie within a normalized 2D space of $[0,1]\times[0,1]$. For each grid $\mathbb{x}_{g}=(\frac{i}{N_{w}},\frac{j}{N_{h}})$, we compute a value $\hat{s}(\mathbb{x}_{g})$ based on the local metric distribution, which is then colored using a continuous colormap. $N_{w}$ and $N_{h}$ denote the number of grids along the x-axis and y-axis, respectively. In this manner, we can generate a contour map depicting continuous landscape of the metric distribution from the discrete projected sample points.
@@ 4.1.1 Kernel Regression Supervised Projection
where the overall regression loss $MSE_{r}$ is a weighted sum of the loss on validation set ($MSE_{vl}$) and that on training set ($MSE_{tr}$). We seek to balance these loss terms to ensure mapping accuracy at both in-sample and out-of-sample positions on the contour, and the weights $w_{1}=1$ and $w_{2}=0.3$ are set empirically. The construction of this weighted train-val loss may differ from common practice but is motivated by our deeper thinking on the problem of connecting projection with kernel regression, which is further explained as follows. Unlike neural network methods, kernel regression is nonparametric, where the parameter like bandwidth is traditionally either precomputed or optimized via cross-validation. Thus, we are facing an interesting problem of how to bridge between parametric projection and nonparametric kernel regression, motivating us to leave a validation set to learn kernel parameters jointly with the neural network. The validation loss improves the generalization of kernel regression for unseen positions. However, relying solely on the validation loss will decrease the local detail of the map. In fact, as shown by additional experiments in Figure 14 in Appendix E, this train-val scheme proves to be essential to ensure robust mapping.
@@ 4.2 Cross-modal Metric Visualization
  • Overlay. The scatterplot view and contour map allow data points to be overlaid onto the contour map, enabling users to simultaneously observe the overall distribution while interacting with individual data points, as shown in Figure 4. To mitigate occlusion, we have implemented sampling techniques, such as random sampling and Poisson disk sampling.
@@ 5 Experiment
Table 1: Quantitative comparison of cross-modal embedding metric visualizations for test (out-of-sample) points on the HPD dataset. Our AKRMap outperforms baseline methods and ablations, achieving the best performance across all four cross-modal embedding metrics.
@@ 5.1 Quantitative Experiments
Baseline methods. We compare AKRMap against three commonly used DR methods: PCA, t-SNE and UMAP, as well as two encoder-based approaches: SAE (Le et al., 2018) and Neuro-Visualizer (Elhamod & Karpatne, 2024). For the encoder-based methods, the decoder is used to estimate grid colors, while the other baseline methods rely on the common Gaussian RBF kernel to estimate the 2D distribution map. Additional details about RBF parameter selection are provided in Appendix C. We use distance threshold to cut off empty areas for traditional DR techniques but keep the full landscape of autoencoder methods because far away positions in traditional DR results are ambiguous and can hardly be mapped back to original high-dimensional space without an inverse decoder. On the other hand, as autoencoder can easily lead to value explosion, we set an upper bound for the map to align with the maximum metric value in the dataset.
@@
Cross-modal Embedding Metrics. We compare AKRMap against baseline methods for visualizing several commonly used cross-modal embedding metrics in T2I generation, including CLIPScore (Hessel et al., 2021), HPSv2 (Wu et al., 2023a), PickScore (Kirstain et al., 2023), and a unimodal embedding metric commonly used in this cross-modal scenario, the Aesthetic Score (Schuhmann et al., 2022).
@@ 5.2.1 Visualizing Large T2I Dataset
In alignment with the results from the quantitative experiments, our method demonstrates superior performance by accurately depicting score distributions with rich local details. In contrast, the baselines suffer from various limitations, such as over-smoothing due to averaging effects, which obscure local structures. For example, PCA introduces pronounced block effects with sharp edges at region boundaries, significantly disrupting the overall smoothness of the visualizations. Similarly, UMAP and t-SNE often struggle with low-value regions being either overshadowed by high values or averaged out, making it challenging to identify clusters with suboptimal performance. Surprisingly, Neuro-Visualizer performs the worst among the baselines for visualizing cross-modal embedding metrics. Except the Aesthetic Score, its results are highly unstable, with excessively large estimations in out-of-sample areas, and the contour maps are riddled with jagged terrains, exhibiting poor smoothness. This highlights the increased challenge of cross-modal embeddings compared to unimodal embeddings. It also further underscores the strength of our AKRMap in producing more accurate and visually coherent representations. Overall, our AKRMap achieves a superior balance between local accuracy and smoothness. Furthermore, the mapping performance of traditional methods can deteriorate significantly if a smaller bandwidth is manually set, as demonstrated by additional results in Appendix C.